A scalloped hammerhead shark swims at a steady speed of 1.5 m/s with its 87-cm-wide head perpendicular to the earth's 50 μT magnetic field.

Part A

What is the magnitude of the emf induced between the two sides of the shark's head? Express your answer using two significant figures. answer must be in mV
Answer #1

here,

speed of the shark , v = 1.5 m/s

width of the head , l = 0.87 m

the earth's magnetic feild , B = 50 uT

B = 5 * 10^-5 T

the magnitude of the emf induced between the two sides of the shark's head , EMF = B * v * L

EMF = 5 * 10^-5 * 1.5 * 0.87

EMF = 6.525 * 10^-5 V

EMF = 0.065 mV

the magnitude of the emf induced between the two sides of the shark's head is 0.065 mV